OpenStudy (anonymous):

Solve each proportion. Leave your answer as a fraction in simplest form. 6/2=4/p (6 over 2 equals 4 over P)

jimthompson5910 (jim_thompson5910):

6/2=4/p 6*p = 2*4 ... cross multiply 6p = 8 p = ??